OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!!!!!!!!!!!!!!!!! Zach borrowed 4,000 from his grandfather to buy a used car so he could travel to and from work. After working for 2 years, he has saved 4,720 to pay back to his grandfather. At what interest rate did Zach borrow the money?

jimthompson5910 (jim_thompson5910):

Interest = Amount Paid back - Amount Borrowed = 4720 - 4000 = 720 Interest = Principle*Rate*Time I = Prt 720 = 4000*r*2 720 = 8000r 720/8000 = r 0.09 = r r = 0.09 So the interest rate was 9%

OpenStudy (anonymous):

THANKS!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

you're welcome